OpenStudy (anonymous):

What value of k makes the equation true? 12 + (–17) = –17 + k k = 12

OpenStudy (jacobciezki):

well what you think

OpenStudy (anonymous):

12

OpenStudy (anonymous):

the answer is right

OpenStudy (jacobciezki):

no

OpenStudy (anonymous):

correct great job you are very smart

OpenStudy (anonymous):

yes huh

OpenStudy (jacobciezki):

wait its not 12 how you get 12

OpenStudy (anonymous):

-17 + 12 is -5

OpenStudy (anonymous):

and 12- 17 is -5 my goodness smh

OpenStudy (jacobciezki):

no

OpenStudy (anonymous):

hey where are you??

OpenStudy (anonymous):

yes it's 12 . 12+(-17)=-17+k or, 12-17=-17+k or, 12=k

OpenStudy (anonymous):

o_o ?>

OpenStudy (anonymous):

plz medal me

OpenStudy (jacobciezki):

why

OpenStudy (anonymous):

see i was right

OpenStudy (anonymous):

i was the first to give the right answer

OpenStudy (jacobciezki):

idc who

OpenStudy (anonymous):

????????????????im confused

OpenStudy (anonymous):

its 12 or -5 :'(

OpenStudy (jacobciezki):

nothing ok your right its 12

OpenStudy (anonymous):

*hugs jacob* ^_^ oaky

OpenStudy (jacobciezki):

thx ^_^

OpenStudy (anonymous):

the answer is 12 but he insisted that it was negative -5

myininaya (myininaya):

Hi. 12+(-17) equals -17+12 by the commutative property of addition that is 12+(-17)=-17+12 comparing this to 12+(-17)=-17+k you should see k is 12 as already mentioned above

OpenStudy (jacobciezki):

/yep she got it

myininaya (myininaya):

Also I hope you guys are being nice to each other.

myininaya (myininaya):

Great.

OpenStudy (jacobciezki):

yours